OpenStudy (anonymous):

differentiate 2sin^2x + sin2x^2

OpenStudy (anonymous):

correction: sin^2(2x) + sin2x^2

OpenStudy (anonymous):

so you're saying \[\sin^2(2x)+\sin(2*x^2)\] or is it \[\sin^2(2x)+\sin((2x)^2)\] ?

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

ok then you can also write it like: \[(\sin(2x))^2+\sin(2x^2)\] and use the chain rule for differentiation in both terms: The derivative of the first term is \[2(\sin(2x))^1*\cos(2x)*2\] The derivative of the second term is \[\cos(2x^2)*4x\] The answer is the two of these derivatives added together. Does that help?

OpenStudy (anonymous):

thanks alot..

OpenStudy (anonymous):

:)